OpenStudy (anonymous):

which of the following are 5th roots of the complex number 32(cos(pi/3) +isin(pi/3))? a. 32(cos(pi/15) +isin(pi/15)) b. 2(cos(7pi/15) +isin(7pi/15)) c. 2(cos(11pi/15) +isin(11pi/15)) d. 2(cos(pi/15) +isin(pi/15)) check all that apply @mustafa2014

OpenStudy (anonymous):

@linn99123 @undeadknight26

undeadknight26 (undeadknight26):

didn't learn this yet so sorry.

OpenStudy (anonymous):

@phi

OpenStudy (phi):

first, you want the 5th root of 32. any idea what that is ?

OpenStudy (phi):

2^5 = 32 so you want a 2 out front the 5th root of cos(x) + i sin(x) is cos(x/5) + i sin(x/5) also, there are 5 roots located at even intervals around the circle. they are spaced every 2 pi/5 or 6 pi/15 radians

OpenStudy (anonymous):

@phi what about the 3rd root of 27 (cos(pi/5) + isin(pi/5))? i know 3 goes out front but i dont know the rest

OpenStudy (phi):

divide the angle by 3 to pi/15 there are 3 roots, space 2pi/3 radians or 10 pi/15 so the roots are at pi/15, pi/15+ 10pi/15 = 11 pi/15 and pi/15+ 20pi/15 = 21 pi/15 (or 7 pi /5 )

OpenStudy (anonymous):

thank you so much!!!!!!!!

OpenStudy (phi):

yw